OpenStudy (anonymous):

x^3=2^(1-x) Solve for x

OpenStudy (anonymous):

\[\Large x^3=2^{1-x}\]

geerky42 (geerky42):

It is unsolvable with only elementary algebra, you would have to go beyond that.

OpenStudy (anonymous):

The answer is 1 but I want to know the intermediate steps

OpenStudy (anonymous):

What are you referring to? @geerky42

geerky42 (geerky42):

I mean to solve it, you would have to use some advanced math So to know that x = 1, you would just have to observe that at x=1, \(x^3 = 1\) and \(2^{1-x} = 2^{1-1} = 2^0 = 1\)

OpenStudy (anonymous):

OK, thanks anyways since it seems that it's really too complicated to even explain, I;m in calculus if that helps xD

geerky42 (geerky42):

yeah, well, you could solve with calculus, probably, but steps is so complicated I cannot comprehend it. You would need to use Lambert W-Function to solve that: http://mathworld.wolfram.com/LambertW-Function.html

OpenStudy (anonymous):

Thanks

geerky42 (geerky42):

welcome